# Cosmic evolution of the Kantowski-Sachs universe in the context of a bulk viscous string in teleparallel gravity

### v4

@@ Abstract
In the present work, we analyzed the Kantowski-Sachs cosmological model and teleparallel gravity, where a bulk viscous fluid containing one-dimensional cosmic strings is the source for the energy--momentum tensor. To obtain the deterministic solution of the field equations, we employed the proportionality condition linking the shear scalar $(\sigma)$ and the expansion scalar $(\theta)$, establishing a relationship between metric potentials. Another approach employed is the Hybrid Expansion Law (HEL). The discussion focuses on the behaviour of the accelerating universe concerning the specific choice of a nonlinear (or power law model) of teleparallel gravity $f(T)=\alpha T+\beta T^{m}$, where $T$ is the torsion scalar, $\alpha$ , and $\beta$ are model parameters and $m$ is restricted to greater than or equal to 2. The effective equation of the state parameter $(\omega_{eff})$ of models I, II and III will support the acceleration behavior of the universe. We observed that the null and weak energy conditions are obeyed but violate the strong energy condition as per the present accelerating expansion. Under specific model parameter constraints, the universe shows a transition from a decelerating to an accelerated phase.
Keywords: Kantowski--Sachs universe, bulk viscous string, hybrid expansion law, $f(T)$ gravity.
August 14, 2024
@@ I Introduction
The structure of this paper is as follows: In section (II) we describe the formulation of $f(T)$ gravity. Section (III) contains a brief review of the Kantowski-Sachs universe model along with the field equations. In section (IV), we describe the field equations and study the three models which are based on the value of a model parameter. This section also contains a comprehensive examination of physical parameters and kinematic parameters. Section (V) contains the energy conditions of models I to III, and in section (VI), we conclude the paper.
@@ II Overview of $f(T)$ gravity
A teleparallel structure that is directly related to the gravitational field generated in space-time by a nontrivial tetrad field. Otherwise, it is possible to define the so-called Weitzenböck connection given a nontrivial tetrad.
@@
where $\rho=\rho_{p}+\lambda$ is the total energy density with particles attached to them, $\rho_{p}$ is the particle energy density, $\lambda$ is the string tension density, $p-3\xi H$ is the bulk viscous pressure, $\xi(t)$ is the coefficient of bulk viscosity, $H$ is Hubble’s parameter, $x^{\nu}$ is a unit space-like vector for cloud string and $u^{\nu}$ denotes the four velocities satisfying the conditions $u^{\nu}u_{\nu}=1=-x^{\nu}x_{\nu}$ and $u_{\nu}x^{\nu}=0$.
@@ III Kantowski-Sachs universe
Various studies have been conducted on Kantowski-Sachs geometries due to their importance [87, 88, 89, 90]. The effects of the cosmological constant on GR have been studied [91, 92], while the case of a perfect fluid obeying the barotropic equation of state was examined by C.B. Collins et al. [88]. The presence of the cosmological constant leads the Kantowski-Sachs universe to become a de Sitter universe in the future [91], making it suitable for describing the preinflationary era [92]. In [88], it was discovered that models with a fluid source exhibit a past asymptotic behavior resulting in a big-bang singularity and a future attractor leading to a large crunch. The effects of the electromagnetic field were explored in [93, 94, 95]. The significance of the Kantowski-Sachs geometry can be observed in inhomogeneous models. For silent universes, this importance is clear [96]. In the case of Szekeres spacetimes, [97] the field equations of Kantowski-Sachs restrict the dynamical variables of the field equation system. This means that there are silent in homogeneous and anisotropic universes that reduce to the anisotropic Kantowski-Sachs geometry when homogenized [98]. Ghosh et al. [99] revisited the anisotropic Kantowski-Sachs model in light of a Lorentzian path integral formalism. A. Paliathanasis [100] performed a detailed analysis of the dynamics by investigating the stationary points and reconstructing the Kantowski-Sachs cosmological model in Weyl integrable spacetime with an ideal gas. G.G. Luciano investigated noninteracting and interacting dark energy scenarios for the Kantowski-Sachs universe [101]. The anisotropic and homogeneous Kantowski-Sachs model, derived from a stiff fluid and cosmological constant was investigated by A. Singh for several potential universe scenarios [102].
@@
The field equation for the Kantowski-Sachs Model equation(12) from equations (7)-(11) in $f(T)$ gravity is given as follows:
@@
The deceleration parameter $(q)$ is given as
@@ IV Solution of field equations
To solve nonlinear differential field equations (16)-(18), the following physically conceivable scenarios are considered [103] as
@@
    where $a_{0}$ and $t_{0}$ denote the scale factor and age of the universe today, respectively, and $n>0$ and $b>0$ are the free parameters. As previously noted, the HEL has been utilized to solve the field equation. Only epoch-based universe evolution can be explained by the power-law and exponential-law cosmologies due to the consistency of the deceleration parameter. The universe does not seem to be shifting from a state of deceleration to acceleration in these cosmologies. Kumar [104] and Akarsu et al. [103] take into consideration the following form of the universe’s scale factor to explain such a transition. The HEL, also known as the generalized form of the scale factor, gives rise to a power-law cosmology for $b=0$ and an exponential law for $n=0$, while the combination of $n>0$ and $b>0$ results in a new cosmology arising from the HEL. By choosing this scale factor, we obtain a time-dependent deceleration parameter (eqn. (35) and fig. (1(a))) that represents inflation and a radiation/matter-dominated era before the dark energy era, followed by a transition from deceleration to acceleration.
@@
Figure 1: These figures show the behaviour of Hubble parameter $H$ and deceleration parameter $q$ at $n=1.5$, $b=0.25$ and $t_{0}=0.5$.
The graphical depictions in Figure (1(a)) showcase the behavior of Hubble parameter $H$. The Hubble parameter exhibits a discernible trend, and its values decrease as cosmic time $t$ progresses. The consistently positive values of $H$ throughout the cosmic evolution unequivocally signify the ongoing expansion of the universe. The Hubble parameter gradually converges toward the smallest positive value.
The deceleration parameter $q$. described by Figure (1(b)), changes over time, showing a transition from the early decelerating phase to the late accelerating phase. This pattern nicely matches what we observe in Type-I supernovae, as $q<0$ for $\frac{t_{0}(\sqrt{n}-n)}{b}<t$. The paper showed that our current universe entered an accelerating phase with a deceleration parameter ranging between $-1\leq q\leq 0$. These results agree well with [109, 110, 111, 112].
@@
In this case, we constrained the model parameters to $\alpha=0$ and $\beta=1$. The field equations are obtained from equations (29)and (30). Equations (16)-(18) are rewritten as
@@
In this case, we consider the values of the model parameters $\alpha=1$ and $\beta=-1$. Rewriting the field equations (16)-(18) with the help of equations (29) and (30), we obtain the new field equations as
@@
To obtain the value of the total energy density $\rho$ of models I to III, we use the field equations (39), (42) and (45). The total energy density for each model is calculated as
@@
The proper energy density $\rho$ exhibits a swift decline from positive values to the same constant, as shown in figures (2(a)), (2(b)), and (2(c)) for models I, II, and III, respectively. The substantial initial values of $\rho$ signify the domination of densities during the early stages, which decreases significantly as time progresses. Importantly, $\rho$ maintains positive values throughout the entire evolution of the universe. This trend underscores the expansive nature of the universe. According to [79], $\rho\rightarrow 0$, for $t\rightarrow\infty$ the universe approaches a flat universe at a late time. As a result, our model is in good agreement with recent observations, and these respondents strongly agree with [115, 116, 117].
The bulk viscous pressure $\bar{p}$ of models I to III are obtained from field equations (38), (41), and (44) as
@@
The examination of figures (3(a)), (3(b)) and (3(c)) reveals a noteworthy trend for models I, II, and III respectively in the behavior of the bulk viscous ${\bar{p}}$ pressure across the cosmic evolution. The parameter ${\bar{p}}$ consistently exhibits a negative trajectory throughout this evolution, beginning with high negative values later on in models II and III $\bar{p}\rightarrow 0$, while model I approaches a small negative value. This characteristic behavior implies a phase of accelerated expansion in the universe. The persistent negativity in bulk viscous pressure signifies a sustained repulsive effect, contributing to the observed cosmic acceleration; these arguments hold similar results to [115, 116, 117, 118, 119].
The effective equation of state ($\omega_{eff}=\rho/\bar{p}$) for each model is calculated from equations (46) to (51) as
@@
According to observational data $\omega_{eff}$ should remain on the negative axis and less than $-1/3$ for Model I. To remain on the negative axis, we choose the negative value of the model parameter $\beta=-1$. We observed that Model I (Fig. (4(a)) $),\omega_{eff}$ decreases from a higher negative value to a smaller negative value. However, for models II and III (Fig. (4(b)),(4(c))), it becomes reversible, i.e., $\omega_{eff}$ increases from a smaller negative to a higher negative value. For models I and II, $\omega_{eff}\leq-1$. In other words, models I and II remain in the quintessence region, but model II approaches $-1$ as time $t$ progresses; hence, it is possible that model II behaves like a standard $\Lambda$CDM. This finding is in good agreement with [117, 120, 121, 82].
@@
We have noted that the string tension density $\lambda$ of models I and II (Figures. (5(a)) and (5(b))) exhibits a positive correlation with time, always maintaining a positive value until it eventually converges to zero at a subsequent point in time. Letelier [71] highlighted the potential occurrence of both positive and negative values for $\lambda$. In model III (Fig. (5(c))), $\lambda$ remains negative throughout the evolution and converges to zero for a late time. For model III where $\lambda<0$, the string phase of the universe ceases to exist, resulting in the presence of an anisotropic fluid composed of particles. Our findings show comparable outcomes to those of [118, 119].
The coefficient of bulk viscosity $\xi$ for this model is obtained from equations (33) and (46)-(48) as
@@
The coefficient of bulk viscosity $\xi$ consistently decreases with time, maintaining a positive value throughout cosmic evolution for each model, as shown in (6(a)), (6(b)), and (6(c)). They ultimately lead to the emergence of inflationary models. This observed behaviour aligns well with the anticipated physical characteristics of $\xi$, and these reports are in good agreement with [122, 123].
@@
Equation (61) implies that the anisotropic parameter $A_{m}$ is time-independent but depends on the values of parameter $k$, as shown in Figure (7(a)). It is observed that the anisotropic parameter has a singularity at $k=-2$, and for $k=1$, it shows isotropic behavior. However, it becomes anisotropic when $k>1$.
The expansion scalar $\theta$ is a positive anti-exponential time-dependent parameter that converges to zero as time $t$ progresses, as shown in Figure (7(b)). Like the expansion scalar $\theta$, the shear scalar $\sigma^{2}$ is also time-dependent and similar to the anisotropic parameter $A_{m}$, it depends on the value of parameter $k$, which is determined from equation (63) and figure (7(c)). Figure (7(c)) shows that for any value of ($k\neq-2$), $\sigma^{2}\rightarrow 0$ as $t\rightarrow\infty$ but for $k>1$, it approaches a positive trajectory and for $k<1$, it approaches a negative trajectory, as shown in Figure (7(c)).
@@ V Energy Conditions
Here, we have described the spatially homogeneous and anisotropic Kantowski-Sachs space-time in the presence of a bulk viscous fluid with one-dimensional cosmic strings in $f(T)$ gravity by modifying general relativity to explain the challenging problem of late time acceleration of the universe. To obtain a definite solution to the highly non-linear field equations (17)--(19) of this theory, we use the special law of variation (26)--(28).
@@
Interestingly, the Hubble parameter $H$ decreases and ultimately approaches a small positive value within the passage of cosmic time. Figure (1(a)) shows that the model’s expansion is faster at the beginning of the universe and subsequently slows down with time. This demonstrates that the universe’s evolution begins at an infinite rate and decreases as it expands. The deceleration parameter $q$ rapidly decreases and asymptotically approaches $-1$, demonstrating a de-Sitter-like expansion at a late time, which explains the acceleration of the universe at a late time. Figure (1(b)) shows that the universe shows the phase transition of current deceleration to late-time acceleration. Hence, for the future, our models show $H>0$ and $q<0$ implying that the model shows the expanding and accelerating nature of the universe. Figure (2) represents the behaviour of the total energy density $\rho$ for models I, II, and III. It is observed that $\rho$ is a positive decreasing function of time, which shows the expansion nature of the universe. The negative behavior of the bulk viscous pressure $\bar{p}$ shown in figure (3) corresponds to cosmic acceleration according to standard cosmology. Every model shows negative trajectories throughout cosmic evolution. Model I approaches a small negative value, while models II and II approach zero for a later time. Therefore, constructed models of the universe show an accelerated expansion phase in the future. One important parameter that characterizes many matter-dominated periods in the development of the universe is the EoS parameter $\omega_{eff}$. Figure (4) depicts the behavior of the EoS parameter against cosmic time $t$. Through an analysis of Model I, (4(a)), we found that the EoS parameter takes on the value $\omega_{eff}\sim-1$ during the early stages of the universe’s history and then entered a quintessence era for later times. This result is consistent with certain EoS constraints that have recently been linked to the inflationary EoS. As time $t$ progresses, model II in Figure (4(b)) behaves like the $\Lambda$CDM model. For Model III, the effective EoS parameter $\omega_{eff}$ figure (4(c)) resembles the dark energy model and behaves like the quintessence model $-1<\omega_{eff}<0$ throughout the cosmic evolution. The string density $\lambda$ is a decreasing function of time and is always greater than zero for models I and II. Figures (5(a)) and (5(b)) show the plots of string tension density versus time $t$ for models I and II. The $\lambda$ remains positive for both models. However, it decreases more sharply with increasing cosmic time. In the early phase of the universe, the string tension density of both models will dominate the dynamics, and later, it approaches zero. However, this becomes a reversible case for Model III. The string energy density $\lambda$ is negative throughout cosmic evolution. Figure (5(c)) shows that at late times, $\lambda$ approaches zero. Finally, we can conclude that for every model at the present epoch, we obtain a string-free model, and the string phase of the universe disappears, i.e., we have an anisotropic fluid of particles. The cosmic fluid dynamics of the universe are significantly affected by the positive decrease and ultimate decrease in the coefficient of bulk viscosity $\xi$ for model II (6(b)), but the coefficient converges to a small positive value for models I and III [Figs. (6(a)) and (6(c)). This decrease in bulk viscosity might lead to a further acceleration of the universe’s expansion, potentially causing it to expand at an even faster rate. As a result, the universe could enter a phase of rapid expansion, similar to inflationary models, which have been proposed to explain the observed accelerated expansion Furthermore, the decrease in viscosity could help prevent initial singularities, consistent with models indicating the beginning of a smooth universe. Additionally, a decrease in bulk viscosity can enhance the role of dark energy or modified gravity in driving accelerated expansion. From figures (7(a)) and (7(c)), we observe that for $k=1$, every model becomes isotropic and shear-free. From the figure (8), we confirmed that for models I to III, the WEC and DEC are valid but violate the SEC. The violation of the SEC has an anti-gravitational effect for which the universe is jerked; thus, our universe exhibits a transition from early decelerating to the present accelerating universe.
@@
Understanding the behavior of the accelerating universe under the influence of teleparallel gravity can provide insights into cosmic evolution. The study of the transition from a decelerated to an accelerated phase in the universe can contribute to our knowledge of the dynamics of the cosmos. Investigating the three different models of $f(T)$ gravity which is a specific form of teleparallel gravity and its impact on the acceleration behaviour of the universe may have implications for cosmological models and theories are the major applications of this constructed model. Further exploration of the impact of the bulk viscousity content in various cosmological scenarios could enhance our understanding of cosmic evolution. Investigating the behavior of bulk viscous pressure and pressure parameters across cosmic evolution may lead to insights into the mechanisms driving cosmic acceleration. Future studies could focus on refining the highly nonlinear form of teleparallel gravity used in the model to improve the accuracy of predictions regarding the accelerating universe. Every constructed model has some loopholes, and the model constraints may limit the generalizability of our findings. The highly nonlinear form of teleparallel gravity used in the model may introduce complexities that could affect the interpretability of the results.

### v6

@@ Abstract
In the present work, we analyzed the Kantowski-Sachs cosmological model and teleparallel gravity, where a bulk viscous fluid containing one-dimensional cosmic strings is the source for the energy--momentum tensor. To obtain the deterministic solution of the field equations, we employed the proportionality condition linking the shear scalar $(\sigma)$ and the expansion scalar $(\theta)$, establishing a relationship between metric potentials. Another approach employed is the hybrid expansion law (HEL). The discussion focuses on the behavior of the accelerating universe concerning the specific choice of a nonlinear (or power law model) of teleparallel gravity $f(T)=\alpha T+\beta T^{m}$, where $T$ is the torsion scalar, $\alpha$ , and $\beta$ are model parameters and $m$ is restricted to greater than or equal to 2. The effective equation of the state parameter $(\omega_{eff})$ of models will support the acceleration of the universe. We observed that the null and weak energy conditions are obeyed but violate the strong energy condition as per the present accelerating expansion. Under specific model parameter constraints, the universe shows a transition from a decelerating to an accelerating phase.
Keywords: Kantowski--Sachs universe; Bulk viscous string; Hybrid expansion law; $f(T)$ gravity.
September 15, 2024
@@ I Introduction
The structure of this paper is as follows: In Sect. (II) we describe the formulation of $f(T)$ gravity. Section (III) contains a brief review of the Kantowski-Sachs universe model along with the field equations. In Sect. (IV), we describe the field equations and study the three models, which are based on the value of a model parameter. This section also contains a comprehensive examination of physical parameters and kinematic parameters. Section (V) contains the energy conditions of models I to III, and in Sect. (VI), we conclude the paper.
@@ II Overview of $f(T)$ gravity
A teleparallel structure which is directly related to the gravitational field generated in space-time by a nontrivial tetrad field. Otherwise, it is possible to define the so-called Weitzenböck connection given as nontrivial tetrad.
@@
where $\rho=\rho_{p}+\lambda$ is the total energy density with particles attached to them, $\rho_{p}$ is the particle energy density, $\lambda$ is the string energy density, $p-3\xi H$ is the bulk viscous pressure, $\xi(t)$ is the coefficient of bulk viscosity, $H$ is Hubble parameter, $x^{\nu}$ is a unit space-like vector for cloud string and $u^{\nu}$ denotes the four velocities satisfying the conditions $u^{\nu}u_{\nu}=1=-x^{\nu}x_{\nu}$ and $u_{\nu}x^{\nu}=0$.
@@ III Kantowski-Sachs universe
Various studies have been conducted on Kantowski-Sachs geometries due to their importance [87, 88, 89, 90]. The effects of the cosmological constant on GR have been studied [91, 92], while the case of a perfect fluid obeying the barotropic equation of state was examined by C.B. Collins et al. [88]. The presence of the cosmological constant leads the Kantowski-Sachs universe to become a de Sitter universe in the future [91], making it suitable for describing the preinflationary era [92]. In [88], it was discovered that models with a fluid source exhibit a past asymptotic behavior resulting in a Big-Bang singularity and a future attractor leading to a large crunch. The effects of the electromagnetic field were explored in [93, 94, 95]. The significance of the Kantowski-Sachs geometry can be observed in homogeneous models. For silent universes, this importance is clear [96]. In the case of Szekeres spacetimes, [97] the field equations of Kantowski-Sachs restrict the dynamical variables of the field equation system this means that there are silent in homogeneous and anisotropic universes that reduce to the anisotropic Kantowski-Sachs geometry when homogenized [98]. Ghosh et al. [99] revisited the anisotropic Kantowski-Sachs model in light of a Lorentzian path integral formalism. A. Paliathanasis [100] performed a detailed analysis of the dynamics by investigating the stationary points and reconstructing the Kantowski-Sachs cosmological model in Weyl integrable spacetime with an ideal gas. G.G. Luciano investigated noninteracting and interacting dark energy scenarios for the Kantowski-Sachs universe [101]. The anisotropic and homogeneous Kantowski-Sachs model, derived from a stiff fluid and cosmological constant was investigated by A. Singh for several potential universe scenarios [102].
@@
The field equations for the Kantowski-Sachs model equation(12) from equations (7)-(11) in $f(T)$ gravity is given as follows:
@@
The deceleration parameter $q$ is given as
@@ IV Solution of field equations
To solve nonlinear differential field equations (16)-(18), the following physically conceivable scenarios are considered [103].
@@
    where $a_{0}$ and $t_{0}$ denote the scale factor and age of the today’s universe, respectively, and $n>0$ and $b>0$ are the free parameters. As previously noted, the HEL has been utilized to solve the field equation. Only epoch-based universe evolution can be explained by the power-law and exponential-law cosmologies due to the consistency of the deceleration parameter. The universe does not seem to be shifting from a state of deceleration to acceleration in these cosmologies. Kumar [104] and Akarsu et al. [103] take into consideration the following form of the universe’s scale factor to explain such a transition. The HEL, also known as the generalized form of the scale factor, gives rise to a power-law cosmology for $b=0$ and an exponential law for $n=0$, while the combination of $n>0$ and $b>0$ results in a new cosmology arising from the HEL. By choosing this scale factor, we obtain a time-dependent deceleration parameter (eqn. (35) and fig. (1(a))) that represents inflation and a radiation/matter-dominated era before the dark energy era, followed by a transition from deceleration to acceleration.
@@
Figure 1: These figures show the behavior of Hubble parameter $H$ and deceleration parameter $q$ at $n=1.5$, $b=0.25$ and $t_{0}=0.5$.
The graphical depictions in Fig. (1(a)) showcase the behavior of Hubble parameter $H$. The Hubble parameter exhibits a discernible trend, and its value decreases as cosmic time $t$ progresses. The consistently positive values of $H$ throughout the cosmic evolution unequivocally signify the ongoing expansion of the universe. The Hubble parameter gradually converges toward the smallest positive value.
The deceleration parameter $q$. displayed in Fig. (1(b)), changes over time, showing a transition from the early decelerating phase to the late accelerating phase. This pattern nicely matches what we observe in Type-I supernovae, as $q<0$ for $\frac{t_{0}(\sqrt{n}-n)}{b}<t$. Hence, the universe has entered an accelerating phase with a deceleration parameter ranging between $-1\leq q\leq 0$. These results agree well with [109, 110, 111, 112].
@@
In this case, we constrained the model parameters to $\alpha=0$ and $\beta=1$. The field equations are obtained from Eqs. (29)and (30) then Eqs. (16)-(18) are rewritten as
@@
In this case, we consider the values of the model parameters $\alpha=1$ and $\beta=-1$. Rewriting the field Eqs. (16)-(18) with the help of Eqs. (29) and (30), we obtain the new field equations as
@@
To obtain the value of $\rho$ for the models I to III, we use the field Eqs. (39), (42) and (45). The total energy density for each model is calculated as
@@
The proper energy density $\rho$ exhibits a swift decline from positive values to the same constant, as shown in Figs. (2(a)), (2(b)), and (2(c)) for models I, II, and III, respectively. The substantial initial values of $\rho$ signify the domination of densities during the early stages, which decreases significantly as time progresses. Importantly, $\rho$ maintains positive values throughout the entire evolution of the universe. This trend underscores the expansive nature of the universe. According to [79], $\rho\rightarrow 0$, for $t\rightarrow\infty$ the universe approaches a flat universe at a late time. As a result, each model is in good agreement with recent observations, and these respondents strongly agree with [115, 116, 117].
The bulk viscous pressure $\bar{p}$ of models I to III are obtained from field Eqs. (38), (41), and (44) as
@@
The examination of Figs (3(a)), (3(b)) and (3(c)) reveals a noteworthy trend for models I, II, and III respectively in the behavior of the bulk viscous ${\bar{p}}$ pressure across the cosmic evolution. The parameter ${\bar{p}}$ consistently exhibits a negative trajectory throughout this evolution, beginning with high negative values later on the models II and III show $\bar{p}\rightarrow 0$, while model I approach a small negative value. This characteristic behavior implies the accelerated expansion of the universe. The persistent negativity in bulk viscous pressure signifies a sustained repulsive effect, contributing to the observed cosmic acceleration; these arguments hold similar results to [115, 116, 117, 118, 119].
The effective equation of state ($\omega_{eff}=\rho/\bar{p}$) for each model is calculated from Eqs. (46) to (51) as
@@
According to, observational data $\omega_{eff}$ should remain on the negative axis and less than $-1/3$ for model I. To remain on the negative axis, we choose the negative value of the model parameter $\beta=-1$. We observed that (Fig. (4(a)) $),\omega_{eff}$ for model I decreases from a higher negative value to a smaller negative value. However, for models II and III (Fig. (4(b)),(4(c))), it becomes reversible, i.e., $\omega_{eff}$ increases from a smaller negative to a higher negative value. For models I and II, $\omega_{eff}\leq-1$. In other words, models I and II remain in the quintessence region, but model II approaches $-1$ as time $t$ progresses; hence, it is possible that model II behaves like a standard $\Lambda$CDM. This finding is in good agreement with [117, 120, 121, 82].
@@
We have noted that the string energy density $\lambda$ of models I and II (Fig. (5(a)) and (5(b))) exhibits a positive correlation with time, always maintaining a positive value until it eventually converges to zero at a subsequent point in time. Letelier [71] highlighted the potential occurrence of both positive and negative values for $\lambda$. In model III (Fig. (5(c))), $\lambda$ remains negative throughout the evolution and converges to zero for a late time. For model III where $\lambda<0$, the string phase of the universe ceases to exist, resulting in the presence of an anisotropic fluid composed of particles. Our findings show comparable outcomes to those of [118, 119].
The coefficient of bulk viscosity $\xi$ for the models are obtained from Eqs. (33) and (46)-(48) as
@@
The coefficient of bulk viscosity $\xi$ consistently decreases with time, maintaining a positive value throughout cosmic evolution for each model, as shown in Figs. (6(a)), (6(b)), and (6(c)). They ultimately lead to the emergence of inflationary models. This observed behavior aligns well with the anticipated physical characteristics of $\xi$, and these reports are in good agreement with [122, 123].
@@
Equation (61) implies that the anisotropic parameter $A_{m}$ is time-independent but depends on the values of the parameter $k$, as shown in Fig. (7(a)). It is observed that the anisotropic parameter has a singularity at $k=-2$, and for $k=1$, it shows isotropic behavior. However, it becomes anisotropic when $k>1$.
The expansion scalar $\theta$ is a positive anti-exponential time-dependent parameter that converges to zero as time $t$ progresses, as shown in Fig. (7(b)). Like the expansion scalar $\theta$, the shear scalar $\sigma^{2}$ is also time-dependent and similar to the anisotropic parameter $A_{m}$, it depends on the value of the parameter $k$, which is determined from Eqn. (63) and Fig. (7(c)). Figure (7(c)) shows that for any value of ($k\neq-2$), $\sigma^{2}\rightarrow 0$ as $t\rightarrow\infty$ but for $k>1$, it approaches a positive trajectory and for $k<1$, it approaches a negative trajectory.
@@ V Energy Conditions
Here, we have described the spatially homogeneous and anisotropic Kantowski-Sachs space-time in the presence of a bulk viscous fluid with one-dimensional cosmic strings in $f(T)$ gravity by modifying general relativity to explain the challenging problem of late time acceleration of the universe. To obtain a definite solution to the highly non-linear field Eqs. (17)--(19) of this theory, we use the special law of variation (26)--(28).
@@
Interestingly, the Hubble parameter $H$ decreases and ultimately approaches a small positive value within the passage of cosmic time. Figure (1(a)) shows that the model’s expansion is faster at the beginning of the universe and subsequently slows down over time. This demonstrates that the universe’s evolution begins at an infinite rate and decreases as it expands. The deceleration parameter $q$ rapidly decreases and asymptotically approaches $-1$, demonstrating a de Sitter-like expansion at a late time, which explains the acceleration of the universe at a late time. Figure (1(b)) displays that the universe shows the phase transition from current deceleration to late-time acceleration. Hence, for the late time, we get $H>0$ and $q<0$ implying that the models show the expanding and accelerating nature of the universe. Figure (2) represents the behavior of the total energy density $\rho$. It is observed that $\rho$ is a positive decreasing function of time, which shows the expansion nature of the universe. The negative behavior of the bulk viscous pressure $\bar{p}$ is shown in Fig. (3) which corresponds to cosmic acceleration according to standard cosmology. Every model shows negative trajectories throughout cosmic evolution. Model I approaches a small negative value, while models II and II approach zero for a later time. Therefore, the constructed models shows an accelerated expansion phase in the future. One important parameter that characterizes many matter-dominated periods in the development of the universe is the EoS parameter $\omega_{eff}$. Figure (4) depicts the behavior of the EoS parameter against cosmic time $t$. Through an analysis of model I, (4(a)), we found that the EoS parameter takes on the value $\omega_{eff}\sim-1$ during the early stage and then entered a quintessence in era for later times. This result is consistent with certain EoS constraints that have recently been linked to the inflationary EoS. As time $t$ progresses, model II (Fig. (4(b))) behaves like the $\Lambda$CDM model. For model III, the effective EoS parameter $\omega_{eff}$ figure (4(c)) resembles the dark energy model and behaves like the quintessence model $-1<\omega_{eff}<0$ throughout the cosmic evolution. The string density $\lambda$ is a decreasing function of time and is always greater than zero for models I and II. Figures (5(a)) and (5(b)) show the plots of string energy density versus time $t$ for models I and II. The $\lambda$ remains positive for both models. However, it decreases more sharply with increasing cosmic time. In the early phase of the universe, the string energy density of both models will dominate the dynamics, and later, it approaches zero. However, this becomes a reversible case for model III. The string energy density $\lambda$ is negative throughout cosmic evolution. Figure (5(c)) shows that, at late times, $\lambda$ approaches zero. Finally, we can conclude that for every model at the present epoch, we obtain a string-free model, and the string phase of the universe disappears, i.e., we have an anisotropic fluid of particles. The cosmic fluid dynamics of the universe are significantly affected by the positive decrease and ultimate decrease in the coefficient of bulk viscosity $\xi$ for model II (6(b)), but the coefficient converges to a small positive value for models I and III [Figs. (6(a)) and (6(c)). This decrease in bulk viscosity might lead to a further acceleration of the universe, potentially causing it to expand at an even faster rate. As a result, the universe could enter a phase of rapid expansion, similar to inflationary models, which have been proposed to explain the observed accelerated expansion Furthermore, the decrease in viscosity could help to prevent the initial singularities, consistent with models indicating the beginning of a smooth universe. Additionally, a decrease in bulk viscosity can enhance the role of dark energy or modified gravity in driving accelerated expansion. From Fig. (7(a)) and (7(c)), we observe that for $k=1$, every model becomes isotropic and shear-free. From Fig. (8), we confirmed that for models I to III, the WECs and DECs are valid but violate the SEC. The violation of the SECs has an anti-gravitational effect for which the universe is jerked; thus, our universe exhibits a transition from early decelerating to the present accelerating phase.
@@
Understanding the behavior of the accelerating universe under the influence of teleparallel gravity can provide insights into cosmic evolution. The study of the transition from a decelerated to an accelerated phase in the universe can contribute to our knowledge of the dynamics of the cosmos. Investigating the three different models of $f(T)$ gravity which is a specific form of teleparallel gravity and its impact on the acceleration behavior of the universe may have implications for cosmological models and theories are the major applications of this constructed model. Further exploration of the impact of the bulk viscousity content in various cosmological scenarios could enhance our understanding of cosmic evolution. Investigating the behavior of bulk viscous pressure and pressure parameters across cosmic evolution may lead to insights into the mechanisms driving cosmic acceleration. Future studies could focus on refining the highly nonlinear form of teleparallel gravity used in the model to improve the accuracy of predictions regarding the accelerating universe. Every constructed model has some loopholes, and the model constraints may limit the generalizability of our findings. The highly nonlinear form of teleparallel gravity used in the model may introduce complexities that could affect the interpretability of the results.